Please help me!!!!
Using this formula find log2 456.



Answer by stanbon(57967)

You probably meant log2 256 as that can be done without a calculator.
Let x=log2 256
Then 2^x=256
2^x=2^8
x=8
Cheers,
Stan H.
If you did mean log2 456 you can do it with a calculator as follows:
log2 456 = log456 / log2
=2.6589.../0.301...=8.832....